OpenStudy (anonymous):

Solve. 8 – 3x – x = 12 x =

OpenStudy (anonymous):

@Hannah_Waller @amorfide

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

combine like terms

OpenStudy (kidrah69):

-3x-x=?

OpenStudy (anonymous):

I belive X = -1? pretty sure @MS15

OpenStudy (amorfide):

8 – 3x – x = 12 first of all combine algebraic terms -3x-x then you want to bring your numbers to the left and your algebraic terms to the right this is because it is always easier to have a positive algebraic expression so you would subtract 12 from both sides, then you would subtract (-3x -x ) from both sides this would give you 8-12= - ( -3x -x ) since the - outside the bracket just multiplies the inside of the bracket by -1 8-12=3x+x you can finish this surely

OpenStudy (amorfide):

@Hannah_Waller please refrain from posting direct answers, and help people understand how to solve them themselves for future reference as it would be more beneficial towards exams